# PACP: Priority-Aware Collaborative Perception for Connected and Autonomous Vehicles

### v1

@@ 1.1 Background
Recent advances in positioning and perception have emerged as pivotal components in numerous cutting-edge applications, most notably in autonomous driving[1, 2, 3, 4]. These systems heavily rely on precise positioning and acute perception capabilities to safely and adeptly navigate complex road environments. Many solutions exist for positioning and perception in CAVs, including inertial navigation systems, high-precision GPS, cameras, and LiDAR[5]. However, any isolated use of them may not be enough to achieve the desired level of perception quality for safe driving. In contrast, the Bird’s Eye View (BEV) stands out as a more holistic approach. By integrating data from multiple sensors and cameras placed around a vehicle, BEV offers a comprehensive, potentially 360-degree view of a vehicle’s surroundings, offering a more contextually rich understanding of its environment[6]. However, most BEV-aided perception designs have predominantly concentrated on single-vehicle systems. Such an approach may not be enough in high-density traffic scenarios, where unobservable blind spots caused by road obstacles or other vehicles remain a significant design challenge. Therefore, collaborative perception has become a promising candidate for autonomous driving. To mitigate the limitations of the single-vehicle systems, we can leverage multiple surrounding CAVs to obtain a more accurate BEV prediction via multi-sensor fusion[7].
To further reduce the risks of blind spots in BEV prediction, collaborative perception is adapted to enable multiple vehicles to share more complementary surrounding information with each other through vehicle-to-vehicle (V2V) communications[8]. This framework intrinsically surmounts several inherent constraints tied to single-agent perception, including occlusion and long-range detection limitation. Similar designs have been observed in a variety of practical scenarios, including communication-assisted autonomous driving within the realm of vehicle-to-everything[9], multiple aerial vehicles for accuracy perception[10], and multiple underwater vehicles deployed in search and rescue operations[11].
@@ 1.2 State-of-the-Art
V2V collaborative perception: V2V collaborative perception combines the sensed data from different CAVs through fusion networks, thereby expanding the perception range of each CAV and mitigating troubling design issues like blind spots. For instance, Chen et al. [19] proposed the early fusion scheme, which fuses raw data from different CAVs, while Wang et al. [14] employed intermediate fusion, fusing intermediate features from various CAVs, and Rawashdeh et al. [20] utilized late fusion, combining detection outputs from different CAVs to accomplish collaborative perception tasks. Although these methods show promising results under ideal conditions, in real-world environments, where the channel conditions are highly variable, directly applying the same fusion methods often results in unsatisfactory outcomes.
@@ 1.3 Our Contributions
  • This is the first paper to leverage a priority-aware collaborative perception framework involving a carefully designed BEV-match mechanism for autonomous driving, which strikes a balance between the communication overhead and perception accuracy.
  • To enhance transmission quality, we employ a two-stage optimization framework based on submodular theory to jointly optimize transmission rates, link connectivity, and compression ratio, addressing challenges of data-intensive transmissions under constrained time-varying channel capacity.
  • We design and implement a deep learning-based adaptive autoencoder integrated into PACP, assisted by fine-tuning mechanisms at roadside units (RSUs). The experimental evaluation reveals that this approach surpasses the state-of-the-art methods, especially in utility value and AP@IoU.
@@ 2.2 Comparison Between Fairness and Priority
To better understand the limitations of fairness-based schemes, we use two typical fairness-based schemes to showcase their resource allocation and BEV prediction in comparison to the priority-aware scheme. Fig. 1 shows the bandwidth and throughput allocation of three schemes in a V2V network. In our scenario, the ego CAV exploits three schemes to combine three nearby CAVs’ camera perception results for BEV prediction, respectively. The whole bandwidth is 200 MHz. Due to the worst channel condition, it is noted that CAV 1 has the lowest priority, thereby its data is discarded by the priority-aware scheme. Fig. 1(b) shows that the compensation of throughput-fairness scheme for weaker channels by allocating more resources results in a decrease in BEV prediction accuracy. Specifically, if a CAV’s channel quality is extremely poor (e.g., CAV 1), the allocated discrete bandwidth might still be insufficient to ensure equalized throughput. In Fig. 2(a), we show the camera perception results from the ego CAV. In Fig. 2(b), it can be observed that the impact of different communication conditions on the dynamic AP@IoU (mobile objective detection) is significant. Additionally, the superiority of the priority-aware scheme becomes evident in Fig. 2(b), where the dynamic AP@IoU rises to 0.685. Therefore, the priority-aware scheme surpasses the fairness-based methods by discerningly choosing which CAVs to collaborate with, leading to improved BEV predictions.
@@ 3 System Model
In this section, we present a V2X-aided collaborative perception system with CAVs, including the system’s structure, channel modeling, and the constraints of computational capacity and energy.
@@ 4.2 Procedure of Obtaining Priority Weights
where $\hat{T}_{10}=T_{10}\times\varTheta_{10}$ remap coordinates from $\text{BEV}_{1}$ to $\text{BEV}_{0}$. Let $\hat{\varGamma}$ be a point of the BEV feature of CAV 1. Furthermore, the transformation of a point $\hat{\varGamma}$ from the BEV perception $\text{BEV}_{1}$ to $\text{BEV}_{0}$ is denoted by $\hat{\varGamma}^{\prime}=\hat{T}_{10}\times\hat{\varGamma}$. $\hat{\varGamma}^{\prime}$ represents the transformed coordinate in $\text{BEV}_{0}$ and the point $\hat{\varGamma}$ is articulated in homogeneous coordinates, i.e., $\hat{\varGamma}=[x,y,1]^{T}$. Inspired by the definition of IoU, we design a BEV-match mechanism, which exploits BEV-match level as the priority weight. Specifically, the priority weight $\mathcal{P}_{10}$ is formulated as follows:
$\mathcal{P}_{10}=\frac{\|[\mathrm{BEV}_{0}\cap\mathrm{BEV}_{1}^{\prime}]_{\pi}-[\mathrm{BEV}_{0}]_{\pi}\|_{2}}{|[\mathrm{BEV}_{0}]_{\pi}|},$ (11)
where $\text{BEV}_{\text{0}}$ denote the ego CAV’s BEV features and $\text{BEV}_{1}^{\prime}=\hat{T}_{10}\times\text{BEV}_{1}$ signifies the transformed BEV, the operation $[\cdot]_{\pi}$ zeros pixels outside the largest bounding rectangle of the fused BEV’s non-zero region, $\|\cdot\|_{2}$ measures the square root of squared pixel values, ||$\cdot$|| ⋅ | sums all pixel values in an image, and the numerator in Eq. (11) quantifies perception difference between data from CAVs 1 and 0, with the denominator providing normalization. The term $\mathcal{P}$ in fact captures two characteristics between the view of the ego CAV and the transformed view of the assisting CAV for data fusion. If two views are very close, the intersection will be similar, the numerator in Eq. (11) will be very small, thus ${\mathcal{P}}_{10}$ will be very small, which implies that the view from CAV 1 will not be helpful to enhance the perception quality, meaning that the view from CAV 1 can be assigned low priority. If the view from CAV 1 is too dissimilar, the intersection will be close to empty, implying that the numerator in Eq. (11) will be close to 1, meaning that the view from CAV 1 may be more helpful and should be assigned higher priority.
@@ 4.3 Utilization Maximization in Collaborative Perception
The performance of collaborative perception relies on the quality of the V2V communication network. Hence, our aim is to maximize the amount of perception data transmitted under the constraints of communication resources, such as power and bandwidth. However, given the varying contributions of nearby CAVs, it is unwise to directly optimize the total throughput of the V2V networks. Instead, we employ the aforementioned priority weight to adjust the resources allocated to each CAV. Moreover, one group of CAVs might have overlapping perceptual regions, so it is intuitive to leverage the set of the union-style perception region $G_{i}$ as the factor of perception performance, which can be given by $G_{i}=\bigcup_{j\in J_{i}}{\mathcal{G}_{j}}$, where ${\mathcal{G}_{j}}$ denotes the perception region by the $j$th CAV and $J_{i}=\{j|s_{ji}=1\}$. Let $\mathcal{I}\left(\mathcal{S}_{i}\right)$ represent the total area covered by the union of perceptual regions of vehicles in set $J_{i}$, we can define
$\mathcal{I}\left(\mathcal{S}_{i}\right)=\mathcal{A}\left(\bigcup_{j\in J_{i}}{\mathcal{G}_{j}}\right),$ (12)
where $\mathcal{A}(X)$ returns the area of a region $X$. Specifically, given the priority weight $\mathcal{P}$, and the link establishment set $\mathcal{S}$, the sub-utility function for the perception quality is formulated as $\mathcal{U}_{\mathrm{r}}=\sum_{i}\sum_{j=1,j\neq i}{{\mathcal{P}}_{ji}s_{ji}d_{ji}}$, which indicates the accuracy and robustness of surrounding perception. While $\mathcal{U}_{\mathrm{p}}=\sum_{i}\mathcal{I}\left(\mathcal{S}_{i}\right)$ denotes the sub-utility function of the perceptual regions of CAVs, which represents the amount of information collected by data fusion. To maximize both utilities at the same time, the weighted utility function is formulated as:
@@
$\mathbf{P}:\quad\!$ $\max\limits_{\mathcal{\mathcal{R}},\mathcal{S},\mathcal{D}}\quad\mathcal{U}_{\mathrm{sum}}(\mathcal{S},\mathcal{D})$ (14)
@@ 5.1 Nonlinear Programming Problem Analysis
$\mathbf{P_{1}}:\quad\!$ $\max\limits_{\mathcal{R},\mathcal{D}}\ \ \mathcal{U}_{\mathrm{sum}}\left(\mathcal{S}^{(n-1)},\mathcal{D}\right)$ (17)
@@
$\mathbf{P}_{1-1}:\quad$ $\max_{\mathbf{U},\mathcal{D}}\,\,\sum_{j=1}^{N}{\sum_{i=1,i\neq j}^{N}{s_{ij}^{(n-1)}\mathcal{P}_{ij}d_{ij}}}$ (18)
@@
$\mathbf{P_{1-2}}:\quad\!$ $\max\limits_{\mathbf{U}}\ \ \sum_{j=1}^{N}\sum_{i=1,i\neq j}^{N}\frac{s_{ij}^{(n-1)}\mathcal{P}_{ij}u_{ij}}{\max\left({r}_{j,\min},\frac{\eta}{e^{L_{ij}}}\right)}$ (20)
@@
$\mathbf{P_{1-2}}$ is a standard linear programming solvable using techniques like the simplex or interior-point methods. If the optimal outcome of $\mathbf{P_{1-2}}$ is $u_{ij}^{(n)}$, the optimal solutions for transmission rate and adaptive compression ratio are $d_{ij}^{(n)}={u_{ij}^{(n)}}\left[{{\max\left({r}_{j,\min},\frac{\eta}{e^{L_{ij}}}\right)}}^{-1}\right]$ and ${r}_{ij}^{(n)}={u_{ij}^{(n)}}/{d_{ij}^{(n)}}$, respectively. Considering that $d_{ij}$ values are at the edges of the feasible region, the optimal solution for $\mathbf{P_{1-2}}$ matches that of $\mathbf{P_{1-1}}$.

### v2

@@ 1.1 Background
Recent advances in positioning and perception have emerged as pivotal components in numerous cutting-edge applications, most notably in autonomous driving[1, 2, 3, 4, 5]. These systems heavily rely on precise positioning and acute perception capabilities to safely and adeptly navigate complex road environments. Many solutions exist for positioning and perception in CAVs, including inertial navigation systems, high-precision GPS, cameras, and LiDAR[6]. However, any isolated use of them may not be enough to achieve the desired level of perception quality for safe driving. In contrast, the Bird’s Eye View (BEV) stands out as a more holistic approach. By integrating data from multiple sensors and cameras placed around a vehicle, BEV offers a comprehensive, potentially 360-degree view of a vehicle’s surroundings, offering a more contextually rich understanding of its environment[7]. However, most BEV-aided perception designs have predominantly concentrated on single-vehicle systems. Such an approach may not be enough in high-density traffic scenarios, where unobservable blind spots caused by road obstacles or other vehicles remain a significant design challenge. Therefore, collaborative perception has become a promising candidate for autonomous driving. To mitigate the limitations of the single-vehicle systems, we can leverage multiple surrounding CAVs to obtain a more accurate BEV prediction via multi-sensor fusion[8].
To further reduce the risks of blind spots in BEV prediction, collaborative perception is adapted to enable multiple vehicles to share more complementary surrounding information with each other through vehicle-to-vehicle (V2V) communications[9]. This framework intrinsically surmounts several inherent constraints tied to single-agent perception, including occlusion and long-range detection limitation. Similar designs have been observed in a variety of practical scenarios, including communication-assisted autonomous driving within the realm of vehicle-to-everything[10], multiple aerial vehicles for accuracy perception[11], and multiple underwater vehicles deployed in search and tracking operations[12, 13, 14].
@@ 1.2 State-of-the-Art
V2V collaborative perception: V2V collaborative perception combines the sensed data from different CAVs through fusion networks, thereby expanding the perception range of each CAV and mitigating troubling design problems like blind spots. For instance, Chen et al. [22] proposed the early fusion scheme, which fuses raw data from different CAVs, Wang et al. [17] employed intermediate fusion, fusing intermediate features from various CAVs, and Rawashdeh et al. [23] utilized late fusion, combining detection outputs from different CAVs to accomplish collaborative perception tasks. Although these methods show promising results under ideal conditions, in real-world environments, where the channel conditions are highly variable, directly applying the same fusion methods often results in unsatisfactory outcomes. Transmitting raw-level sensing data has distinct advantages: it supports multiple downstream tasks, enhances data fusion accuracy, and ensures future-proof flexibility for evolving autonomous driving technologies. Therefore, we can combine the benefits of early fusion and intermediate fusion by utilizing adaptive compression for raw data transmission.
@@
Priority-Aware Perception Schemes: Priority-aware perception schemes have been advanced significantly, yet they have also encountered limitations in dynamic environments. Liao et al.’s model uses an attention mechanism for trajectory prediction, effectively assigning dynamic weights but overlooking communication costs and information redundancy[29]. Similarly, Wen et al.’s queue-based traffic scheduling enhances throughput but struggles with real-time data synchronization, which is critical for systems like autonomous vehicles[30]. Additionally, studies on edge-assisted visual simultaneous localization and mapping (SLAM) introduce a task scheduler that improves mapping precision but fails to account for the crucial role of channel quality in perception accuracy and throughput[31].
Our framework mitigates these issues by introducing an adaptive mechanism that adjusts priorities based on real-time channel quality and analytics. This approach not only curtails unnecessary data processing and communication but also enhances raw-level sensing data fusion and system responsiveness, thereby addressing the core challenges identified in previous studies.
@@ 1.3 Our Contributions
  • We introduce the first-ever implementation of a priority-aware collaborative perception framework that incorporates a novel BEV-match mechanism for autonomous driving. This mechanism uniquely balances communication overhead with enhanced perception accuracy, directly addressing the inefficiencies found in prior works.
  • Our two-stage optimization framework is the first to apply submodular theory in this context, allowing the joint optimization of transmission rates, link connectivity, and compression ratio. This innovation is particularly adept at overcoming the challenges of data-intensive transmissions under dynamic and constrained channel capacities.
  • We have integrated a deep learning-based adaptive autoencoder into PACP, supported by a new fine-tuning mechanism at roadside units (RSUs). The experimental evaluation reveals that this approach surpasses the state-of-the-art methods, especially in utility value and AP@IoU.
@@ 2.2 Comparison Between Fairness and Priority
To better understand the limitations of fairness-based schemes, we use two typical fairness-based schemes to showcase their resource allocation and BEV prediction in comparison to the priority-aware scheme. Fig. 1 shows the bandwidth and throughput allocation of three schemes in a V2V network. In our scenario, the ego CAV exploits three schemes to combine three nearby CAVs’ camera perception results for BEV prediction, respectively. The whole bandwidth is 200 MHz.
In the subchannel-fairness scheme (Fig. 1(a)), each CAV is allocated an equal amount of spectrum resources, which can lead to inefficient resource utilization, especially when some CAVs have poor channel conditions. The throughput-fairness scheme (Fig. 1(b)) aims to equalize the transmission rates among CAVs by compensating for weaker channels with more resources. However, this approach can still result in suboptimal BEV prediction accuracy. Specifically, if a CAV’s channel quality is extremely poor (e.g., CAV 1), the allocated discrete bandwidth might still be insufficient to ensure equalized throughput, leading to reduced perception performance.
In contrast, the priority-aware perception scheme (Fig. 1(c)) dynamically adjusts the priority weights of each CAV based on channel resources, perception accuracy, and perception coverage. Due to the worst channel condition, it is noted that CAV 1 has the lowest priority, thereby its data is discarded by the priority-aware scheme. This approach ensures that the most critical and high-quality data from the CAVs with the best channel conditions and the closest proximity to the ego vehicle are transmitted with minimum loss and latency. This selective collaboration leads to a significant improvement in BEV predictions. Fig. 2(b) shows the impact of different communication conditions on the dynamic AP@IoU (mobile object detection). The superiority of the priority-aware scheme becomes evident, as the dynamic AP@IoU rises to 0.685. Therefore, the priority-aware scheme surpasses the fairness-based methods by discerningly choosing which CAVs to collaborate with, leading to improved BEV predictions.
@@ 3 System Model
TABLE II: Summary of Key Notations
Notation Definition
$K$ The number of orthogonal sub-channels
$N$ The total number of nodes or vehicles in the network
$C_{ij}$ The channel capacity between the $i$th transmitter and the $j$th receiver
$\mathcal{D}$ The matrix of transmission rates, $\mathcal{D}=\{d_{ij}\}_{N\times N}$
$r_{ij}$ The adaptive compression ratio for data transmission between CAVs $i$ and $j$
$E_{ij}^{t}$ The energy consumption bounds for data transmission between CAVs $i$ and $j$
$E_{j}^{c}$ The energy consumption bounds for computation at CAV $j$
$\mathcal{P}_{ij}$ The priority weight between the ego vehicle $j$ and nearby CAV $i$
$\mathcal{I}(\mathcal{S}_{i})$ The total area covered by the union of perceptual regions of vehicles
$\mathcal{U}_{\mathrm{sum}}$ The weighted utility function for the network
$\mathcal{U}_{\mathrm{r}}$ The sub-utility function for perception quality
$\mathcal{U}_{\mathrm{p}}$ The sub-utility function for the perceptual regions
$\omega_{1},\omega_{2}$ Weights for perception quality and perceptual region, respectively
In this section, we present a V2X-aided collaborative perception system with CAVs, including the system’s structure, channel modeling, the constraints of computational capacity and energy, and background information on IoU. The key notations are listed in the Table II.
@@ 3.2 Channel Modeling and System Constraints
3.3 The Background Information on IoU
Intersection over Union (IoU) is an effective metric in computer vision, especially in object detection, used to quantify the overlap between two areas. Though IoU fundamentally informs the design of our priority weight system in subsequent sections, it is directly applied as a validation tool in our experimental setup to assess the accuracy of perception overlaps among Connected Autonomous Vehicles (CAVs).
(a) The intersection of $G$ and $P$.
(b) The union of $G$ and $P$.
Figure 4: Illustration of Intersection over Union (IoU) using a ground truth bounding box $G$ and a predicted bounding box $P$.
As shown in Fig. 4, IoU is calculated between a ground truth bounding box $G$ and a predicted bounding box $P$, which represent the actual and camera-based perceived locations, respectively. The coordinates for $G$ ($(x_{0}^{G},y_{0}^{G})$ and $(x_{1}^{G},y_{1}^{G})$) and $P$ ($(x_{0}^{P},y_{0}^{P})$ and $(x_{1}^{P},y_{1}^{P})$) denote their respective top-left and bottom-right corners.
The respective areas of $G$ and $P$ are defined as follows:
$S_{G}$ $=|(x_{0}^{G}-x_{1}^{G})\times(y_{0}^{G}-y_{1}^{G})|,$ (10)
$S_{P}$ $=|(x_{0}^{P}-x_{1}^{P})\times(y_{0}^{P}-y_{1}^{P})|.$
The area of overlap, critical for the IoU calculation, is derived from the extremities of the overlapping coordinates:
$S_{G\cap P}=\left|(x_{0}^{I}-x_{1}^{I})\times(y_{0}^{I}-y_{1}^{I})\right|$ (11)
The union area, combining both bounding boxes minus their intersection, is given by:
$S_{G\cup P}=S_{G}+S_{P}-S_{G\cap P}$ (12)
IoU is thus calculated as:
$\mathrm{IoU}(G,P)=\frac{S_{G\cap P}}{S_{G\cup P}}=\frac{\left|(x_{0}^{I}-x_{1}^{I})\times(y_{0}^{I}-y_{1}^{I})\right|}{S_{G}+S_{P}-S_{G\cap P}}.$ (13)
The performance of IoU is heavily influenced by the quality of the communication channel. Superior channel quality enhances the transmission of camera data, leading to more precise and synchronized data fusion among CAVs. This synchronization reduces latency in data exchanges, directly improving the alignment of perceived and actual locations, thus boosting IoU performance. Conversely, poor channel conditions can decrease overlap accuracy between predicted and ground truth boxes, thereby reducing IoU scores. This emphasizes the critical role of robust channel quality in achieving optimal perception system performance in CAV networks.
@@ 4.2 Procedure of Obtaining Priority Weights
where $\hat{T}_{10}=T_{10}\times\varTheta_{10}$ remap coordinates from $\text{BEV}_{1}$ to $\text{BEV}_{0}$. Let $\hat{\varGamma}$ be a point of the BEV feature of CAV 1. Furthermore, the transformation of a point $\hat{\varGamma}$ from the BEV perception $\text{BEV}_{1}$ to $\text{BEV}_{0}$ is denoted by $\hat{\varGamma}^{\prime}=\hat{T}_{10}\times\hat{\varGamma}$. $\hat{\varGamma}^{\prime}$ represents the transformed coordinate in $\text{BEV}_{0}$ and the point $\hat{\varGamma}$ is articulated in homogeneous coordinates, i.e., $\hat{\varGamma}=[x,y,1]^{T}$.
Inspired by IoU, the metric for perceptual quality depends on the intersection of the ground truth and other predicted results. As we are mainly concerned with the impact of an unstable channel on perceptual quality, we can assume that the ego vehicle (CAV 0 in Fig. 6) can obtain highly accurate locations of nearby objects, which can serve as the ground truth. For each nearby CAV within communication range, CAV 0 can calculate priority weights only using overlapping perceptual objects, i.e., we only calculate priority weights based on BEV features (boxes) about objects $\pi_{0},\pi_{1},\pi_{2},$ and $\pi_{3}$. Specifically, the priority weight $\mathcal{P}_{10}$ is formulated as follows:
$\mathcal{P}_{10}=\frac{\|[\mathrm{BEV}_{0}\cap\mathrm{BEV}_{1}^{\prime}]_{\pi}\|_{2}}{|[\mathrm{BEV}_{0}]_{\pi}|},$ (15)
where $\text{BEV}_{\text{0}}$ denote the ego CAV’s BEV features and $\text{BEV}_{1}^{\prime}=\hat{T}_{10}\times\text{BEV}_{1}$ signifies the transformed BEV, the operation $[\cdot]_{\pi}$ zeroes out pixels outside the overlapping perceptual objects of the fused BEV’s non-zero region, $\|\cdot\|_{2}$ measures the square root of squared pixel values, ||$\cdot$|| ⋅ | sums all pixel values in an image, and the numerator in Eq. (15) quantifies perception difference between data from CAVs 1 and 0, with the denominator providing normalization. The term $\mathcal{P}$ in fact captures two characteristics between the view of the ego CAV and the transformed view of the assisting CAV for data fusion. If two views are very close, the intersection will be similar, the numerator in Eq. (15) will be very high, thus ${\mathcal{P}}_{10}$ will be very high, which implies that the view from CAV 1 will accurate enough to enhance the perception quality, meaning that the view from CAV 1 can be assigned high priority. If the view from CAV 1 is too dissimilar, the intersection will be close to empty, implying that the numerator in Eq. (15) will be close to 0, meaning that the view from CAV 1 may be more misleading and should be assigned lower priority.
Figure 6: The example of calculating priority weight $\mathcal{P}_{10}$.
Moreover, our system incorporates a gate mechanism that effectively filters out data from CAVs whose priority weights fall below a certain threshold. This gating process prevents low-quality data from influencing the overall perception quality, thereby maintaining the integrity and accuracy of the collaborative perception system despite the inherent instabilities in vehicular network conditions.
@@ 4.3 Utilization Maximization in Collaborative Perception
The performance of collaborative perception relies on the quality of the V2V communication network. Hence, our aim is to maximize the amount of perception data transmitted under the constraints of communication resources, such as power and bandwidth. However, given the varying contributions of nearby CAVs, it is unwise to directly optimize the total throughput of the V2V networks. Instead, we employ the aforementioned priority weight to adjust the resources allocated to each CAV. Moreover, one group of CAVs might have overlapping perceptual regions. To handle this, we leverage the union-style perception region set $G_{i}$, defined as:$G_{i}=\bigcup_{j\in J_{i}}\mathcal{G}_{j}$, where $\mathcal{G}_{j}$ represents the perception region of the $j$th CAV, and $J_{i}=\{j|s_{ji}=1\}$. Here, $\mathcal{I}(\mathcal{S}_{i})$ represents the total area covered by the union of perceptual regions of vehicles in set $J_{i}$, which is mathematically expressed as:
$\mathcal{I}(\mathcal{S}_{i})=\mathcal{A}\left(\bigcup_{j\in J_{i}}\mathcal{G}_{j}\right)$ (16)
where $\mathcal{A}(X)$ denotes the area of a region $X$. The utilization of the union of regions inherently introduces a submodular property in the coverage function. Submodularity [^2] The definition of submodularity can be found in Definition 3. in this context means that adding an additional CAV to a group of vehicles with significantly overlapping regions provides diminishing marginal gains in covered area. This property naturally discourages the system from including excessive overlaps in the set of actively transmitting vehicles. Specifically, we exploit this submodular property by formulating the sub-utility function for perception quality as $\mathcal{U}_{\mathrm{r}}=\sum_{i}\sum_{j=1,j\neq i}\mathcal{P}_{ji}s_{ji}d_{ji}$, which reflects the accuracy and robustness of surrounding perception. Concurrently, the sub-utility function for the perceptual regions of CAVs is defined as $\mathcal{U}_{\mathrm{p}}=\sum_{i}\mathcal{I}(\mathcal{S}_{i})$, which represents the amount of information collected by data fusion. To maximize both utilities at the same time, the weighted utility function is formulated as:
@@
$\mathbf{P}:$ $\max\limits_{\mathcal{\mathcal{R}},\mathcal{S},\mathcal{D}}\quad\mathcal{U}_{\mathrm{sum}}(\mathcal{S},\mathcal{D})$ (18)
@@ 5.1 Nonlinear Programming Problem Analysis
$\mathbf{P_{1}}:$ $\max\limits_{\mathcal{R},\mathcal{D}}\ \ \mathcal{U}_{\mathrm{sum}}\left(\mathcal{S}^{(n-1)},\mathcal{D}\right)$ (21)
@@
$\mathbf{P}_{1-1}:$ $\max_{\mathbf{U},\mathcal{D}}\,\,\sum_{j=1}^{N}{\sum_{i=1,i\neq j}^{N}{s_{ij}^{(n-1)}\mathcal{P}_{ij}d_{ij}}}$ (22)
@@
$\mathbf{P_{1-2}}:$ $\max\limits_{\mathbf{U}}\ \ \sum_{j=1}^{N}\sum_{i=1,i\neq j}^{N}\frac{s_{ij}^{(n-1)}\mathcal{P}_{ij}u_{ij}}{\max\left({r}_{j,\min},\frac{\eta}{e^{L_{ij}}}\right)}$ (24)
@@
$u_{ij}=0\ \text{if}\ s_{ij}^{(n-1)}=0.$
$\mathbf{P_{1-2}}$ is a standard linear programming solvable using techniques like the simplex or interior-point methods. If the optimal outcome of $\mathbf{P_{1-2}}$ is $u_{ij}^{(n)}$, the optimal solutions for transmission rate and adaptive compression ratio are $d_{ij}^{(n)}={u_{ij}^{(n)}}\left[{{\max\left({r}_{j,\min},\frac{\eta}{e^{L_{ij}}}\right)}}^{-1}\right]$ and ${r}_{ij}^{(n)}={u_{ij}^{(n)}}/{d_{ij}^{(n)}}$, respectively. Considering that $d_{ij}$ values are at the edges of the feasible region, the optimal solution for $\mathbf{P_{1-2}}$ matches that of $\mathbf{P_{1-1}}$. It is noted that $u_{ij}=0\ \text{if}\ s_{ij}^{(n-1)}=0$ guarantees that $u_{ij}$ is explicitly set to zero whenever $s_{ij}^{(n-1)}=0$.